OpenStudy (anonymous):

Solve the logarithmic equation. log 4 x=log 3 + log(x+3)

OpenStudy (asnaseer):

use this log rule to first simplify the right-hand-side:\[\log(a)+\log(b)=\log(ab)\]

OpenStudy (anonymous):

log(3)+log(4x)=log(x+3)??

OpenStudy (asnaseer):

where did you get that from?

OpenStudy (asnaseer):

use the rule I gave to first simplify this:\[\log 3 + \log(x+3)\]

OpenStudy (ksaimouli):